OpenStudy (anonymous):

Evaluate the double integral from 0 to 1 and x^2 to 1 of x^3*sin(y^3) dydx by reversing the order of integration.

OpenStudy (fibonaccichick666):

so we have \[\int^{1}_{0}\int^{1}_{x^2}x^3sin(y^3)dydx\]correct?

OpenStudy (fibonaccichick666):

are you familiar with Fubini's theorem?

OpenStudy (anonymous):

I know how to solve the integral, I just don't know how to set up the integral with the correct limits.

OpenStudy (fibonaccichick666):

Ok so what we want to do is figure out what our bounds on y actually are

OpenStudy (anonymous):

How so? What's the fastest way to find out without graphing?

OpenStudy (fibonaccichick666):

well our second integration in this case is the dx, so we know that our x is constrained as such: 0≤x≤1

OpenStudy (fibonaccichick666):

so what is the smallest x^2 can be?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

I mean, 0?

OpenStudy (fibonaccichick666):

yep why?

OpenStudy (anonymous):

Because I plugged in 0 in x^2.

OpenStudy (fibonaccichick666):

ok good so other than graphing, which is useful when you are eprmitted a calculator, but too time consuming when you do not, you have to use logic. If I show you 0≤x≤1 and x^2≤y≤1 can you just observe and see what the bounds are by thinking?

OpenStudy (anonymous):

Wait a minute, let me work it out.

OpenStudy (anonymous):

The new limits would be 0<=y<=1, x^2<=x<=1, right?

OpenStudy (fibonaccichick666):

not quite you can't have 2 x's in the same inequality try again

OpenStudy (fibonaccichick666):

focus on the second inequality I typed. Could you maybe take the square root of each term?

OpenStudy (anonymous):

So it'd be x<=sqrt(x)<=1, right?

OpenStudy (fibonaccichick666):

no remember you cannot have two x's in the same inequality. why? because then you are left with a variable post integration that you can't do anything with.

OpenStudy (fibonaccichick666):

maybe thinking of it as we want the bounds on x in terms of y will help

OpenStudy (anonymous):

So it's sqrt(y)<=x<=1, right?

OpenStudy (fibonaccichick666):

not quite look at this inequality and tell me if √y is ≤ or ≥ x : x^2≤y≤1